# Conditioning Diffusions Using Malliavin Calculus

### v1

@@ 3.3 Summary of Choices
BEL optimal: The first choice is setting $\alpha$ as in LABEL:{lemma:variance-bel-bm}. Although this only gives optimal variance in the specific case of Brownian motion, it still may do well in other problem settings, especially those similar to Brownian motions.
@@
Reparametrisation: Corresponding to the discussion on the reparametrisation trick in Lemma 3.2, we set $\alpha_{t}=J_{t\mid s}M_{t}$, with $M_{t}=\frac{T-t}{T}\operatorname{Id}$.
@@ 4.1.2 Brownian Motion
$\nabla\log p(X_{1}=x\mid X_{t}=x_{t})=-\frac{x_{t}-x}{1-t}.$ (29)
This allows us to compare our methods against the true bridges. The results are presented in Figure 8. We observe that the reparametrisation trick algorithm performs well for the one-dimensional Brownian motion, while the BEL average achieves the best performance in $10$ dimensions. For an explanation of the metrics see Section C.1.2
@@ 4.1.3 Double-Well
To obtain a ground truth for this example, we numerically estimate $f(x)=p(X_{1}=-1\mid X_{s}=x)$ and compute its logarithmic gradient. We then compare the paths of the process under the true drift with our estimates in Figure 4.
@@
Figure 5: The first and second coordinates of a 10-dimensional process sampled from the SDE in 30 are shown in (a). The process is conditioned on transitioning from $1$ to $-1$ in the first coordinate while remaining at $1$ in all others. Panels (b), (c), and (d) depict the paths generated by the BEL-first, BEL-average, and BEL-optimal algorithms, respectively. All plots include underlying contour lines representing the level sets of the potential.
@@ 4.2 Shape Processes
Our findings are given in Table 1 and show that BEL average outperforms competing methods. The next best method is Heng et al. (2022), which can be interpreted as first applying BEL to the time-reversed process, approximating the transition densities via Gaussians (24a), and using a fixed start point instead of amortisation. However, since their approach involves time-reversal, they do not take gradients of the drift or diffusion terms. BEL average not only uses this gradient information, but non-local information about the score.
@@ 4.3 Image Experiments
Next, we condition the resulting SDE to produce images with a specified upper-left corner. Importantly, this conditioning does not require adding artificial noise to the observation. More specifically, the law of $X_{T}$ is by design the data distribution. Then we condition on $Y=G(X_{T})$ where G selects the upper left corner of the image. Once trained we can therefore choose an arbitrary upper left corner $y$ and sample from images matching this, using the learned additional drift $u(X_{t},y)$. Note we only need train once and then can sample using any corner. We demonstrate the result by conditioning on the upper-left corner of a jacket image in Figure 6.
@@ 5 Conclusion
In this work, we have introduced a novel class of loss functions for estimating the drift of a conditioned diffusion process. Our approach leverages Malliavin calculus and integration by parts to handle singular losses, enabling a more stable and accurate estimation. This framework not only leads to a generalized Tweedie’s formula but also unifies several existing methods under a common theoretical perspective. Empirically, we demonstrated that our loss functions outperform competing algorithms in the simulation of shape processes, and scales to high dimensions on image experiments. Future work includes extending our methodology to a broader range of applications and further investigating the role of $\alpha$, either through analytical techniques or by optimising it.
Table 2: Dimension 1
Loss MV Dist
BEL average $8.1\times 10^{-1}$ $1.6\times 10^{-1}$
BEL first $2.9\times 10^{-1}$ $8.4\times 10^{-2}$
BEL last $1.1\times 10^{0}$ $1.1\times 10^{0}$
BEL optimal $8.3\times 10^{-1}$ $1.9\times 10^{-1}$
Reparametrization Trick $9.2\times 10^{-1}$ $2.1\times 10^{-1}$
Table 3: Dimension 10
Loss MV Dist
BEL average $13.4\times 10^{-}$ 1 $13.0\times 10^{-}$ 1
BEL first $12.1\times 10^{-}$ 1 $13.3\times 10^{-}$ 1
BEL last $15.9\times 10^{-}$ 1 $1.1\times 10^{0}$
BEL optimal $13.1\times 10^{-}$ 1 $13.0\times 10^{-}$ 1
Reparametrization Trick $15.5\times 10^{-}$ 1 $15.2\times 10^{-}$ 1
@@ Acknowledgments
This work is partially supported by the Novo Nordisk Foundation grant NNF18OC0052000, a research grant (VIL40582) from VILLUM FONDEN, and UCPH Data+ Strategy 2023 funds for interdisciplinary research. GD and JP were supported by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
@@ Appendix A Adjoint SDE for Calculating $\mathcal{S}_{s}$
making use of the semigroup property $J_{t|s}=J_{t|s+\delta t}J_{s+\delta t|s}$.
@@
where $\tau$ can be any number in $\tau\in[s,s+\delta t]$. For $\tau=s$, we have $J_{s\mid s}=\text{Id}$. The term $J_{s+\delta t|s}$ can be approximated using an Euler-Maryuama step on 7, leading to $J_{s+\delta t\mid s}\approx\text{Id}+\delta t\nabla b_{s}(X_{s})+\nabla\sigma_{s}(X_{s})\delta B_{s}$.
@@ Appendix B Denoising Score Matching and Tweedies Formula
Here we show how our results can be applied to derive a formula for the score $\nabla\log p_{t}(x)$ of a diffusion process. This could be done by applying Proposition 2.3 to a trivial conditioning $Y=0$ and treating the reverse dynamics of the SDE (6). However, it is instructive to quickly rederive the analogous equation in this simplified setting, which we do in this section. We will see that it generalises Tweedies formula.
@@
and denote by $p_{t}$ the density of $X_{t}$. Then we have the following representation:
@@
$\nabla\log p_{t}(x)=\mathbb{E}\left[\int_{0}^{t}(\sigma(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}|X_{t}=x\right].$
Here $\alpha_{s}^{\prime}$ is a function which satisfies $\int_{0}^{t}\alpha_{s}^{\prime}\,\mathrm{d}s=1$, and $J_{t|s}$ is the Jacobian which follows the flow (7).
@@
$\mathcal{L}(u)=\mathbb{E}\left[\|u(X_{t})-\int_{0}^{t}(\sigma(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\|^{2}\right]$ (39)
has a unique minimiser given by $u(x)=\nabla\log p_{t}(x)$.
@@
$D_{s}\varphi(X_{t})=\nabla\varphi(X_{t})J_{t|s}\sigma_{s}(X_{s}),$
@@
$\nabla\varphi(X_{t})=D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1},$
@@
$\nabla\varphi(X_{t})=\int_{0}^{t}D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1}\alpha_{s}^{\prime}\,\mathrm{d}s$
@@
$\mathbb{E}[\nabla\varphi(X_{t})]=\mathbb{E}\left[\int_{0}^{t}D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1}\alpha_{s}^{\prime}\,\mathrm{d}s\right]=\mathbb{E}\left[\varphi(X_{t})\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\right].$
@@
$=\frac{1}{p_{t}(x)}\mathbb{E}\left[\delta_{x}(X_{t})\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\right]=\mathbb{E}\left[\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}|X_{t}=x\right].$
@@
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}\left[X_{t}-e^{-t/2}X_{0}|X_{t}=x\right],$ (41)
and, in particular, the loss (39) is simplifies to the well-known denoising score matching loss:
@@
$\nabla\log p_{t}(x)=\mathbb{E}\left[\int_{0}^{t}(\sigma(X_{s})^{-1}J_{t|s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}|X_{t}=x\right]=\mathbb{E}\left[\int_{0}^{t}e^{\frac{t-s}{2}}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}|X_{t}=x\right].$
@@
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}\left[\int_{0}^{t}e^{-\frac{t-s}{2}}\mathrm{d}B_{s}|X_{t}=x\right].$ (43)
@@
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}[X_{t}-e^{-t/2}X_{0}|X_{t}=x].$
@@ C.1.3 Results
Here we provide the tables with the results for our controlled environment experiments for Brownian motion Figure 8.
@@
Figure 8: Performance of various algorithms for conditioning a Brownian motion (see Section 4.1.2).
@@ C.2 Shape Processes
Figure 9: We plot one trajectory from the unconditioned stochastic shape process, started at a circle of radius 1.
@@ Appendix D Proofs
$\nabla\log p_{T\mid s}(x_{T}\mid x_{s})$ $=~{}\frac{1}{p_{T\mid s}(x_{T}\mid x_{s})}\nabla_{x_{s}}\mathbb{E}[\delta_{x_{T}}(X_{T})\mid X_{s}=x_{s}]$
$=~{}\mathbb{E}\left[\delta_{x_{T}}(X_{T})\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid X_{s}=x_{s}\right]\frac{A_{T\mid s}^{-1}}{p_{T\mid t}(X_{T}\mid x_{s})}$
@@
The final result of Proposition 2.3 follows transposing $\alpha$ and transposing the last equation; this is only a cosmetical change so that the result is more in line with algorithmic implementations.
Now, using Proposition D.2 for $t=T$ we get that
$\nabla\log p(Y=y|X_{s}=x_{s})$ $=\mathbb{E}[\nabla_{X_{t}}\log p_{t}(X_{T}|X_{s})|Y,X_{s}]$
$=\mathbb{E}\left[\mathbb{E}\left[\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid X_{s},X_{T}\right]|Y,X_{s}\right]$
$=\mathbb{E}\left[\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}|Y,X_{s}\right]$
Proposition D.2.
Let $s\geqslant t$, then we have that
$\nabla_{X_{t}}\log p(Y|X_{s})=\mathbb{E}[\nabla_{X_{t}}\log p_{t}(X_{t}|X_{s})|Y,X_{s}].$
Proof.
We have
$\nabla_{X_{t}}\log p(Y|X_{t})$
$=$ $\frac{1}{p(Y|X_{t})}\nabla_{X_{t}}\int q(Y|X_{T}=x_{T})p(X_{T}=x_{T}|X_{t})\mathrm{d}x_{T}$
$=$ $\frac{1}{p(Y|X_{t})}\nabla_{X_{t}}\int q(Y|X_{T}=x_{T})p(X_{T}=x_{T}|X_{s}=x_{s})p(X_{s}=x_{s}|X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$=$ $\frac{1}{p(Y|X_{t})}\int q(Y|X_{T}=x_{T})p(X_{T}=x_{T}|X_{s}=x_{s})p(X_{s}=x_{s}|X_{t})\nabla_{X_{t}}\log p(X_{s}=x_{s}|X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$=$ $\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}|X_{t})|X_{t},Y].$

### v2

@@ 3.3 Summary of Choices
BEL optimal: The first choice is setting $\alpha_{s}$ as in LABEL:{lemma:variance-bel-bm}. Although this only gives optimal variance in the specific case of Brownian motion, it still may do well in other problem settings, especially those similar to Brownian motions.
@@
Reparametrisation: Corresponding to the discussion on the reparametrisation trick in Lemma 3.2, we set $\alpha_{s}=J_{s\mid t}M_{s}$, with $M_{s}=\frac{T-s}{T}\operatorname{Id}$.
@@ 4.1.2 Brownian Motion
$\nabla\log p_{1\mid t}(X_{1}=x\mid X_{t}=x_{t})=-\frac{x_{t}-x}{1-t}.$ (29)
This allows us to compare our methods against the true bridges. The results are presented in Table 4. We observe that the reparametrisation trick algorithm performs well for the one-dimensional Brownian motion, while the BEL average achieves the best performance in $10$ dimensions. For an explanation of the metrics see Section C.1.2
@@ 4.1.3 Double-Well
To obtain a ground truth for this example, we numerically estimate $f(x)=p_{1\mid s}(X_{1}=-1\mid X_{s}=x)$ and compute its logarithmic gradient. We then compare the paths of the process under the true drift with our estimates in Figure 4.
@@
Figure 5: The first and second coordinates of a 10-dimensional process sampled from the SDE in 30 are shown in (a). The process is conditioned on transitioning from $1$ to $-1$ in the first coordinate while remaining at $1$ in all others. Panels (b), (c), and (d) depict the paths generated by the BEL-first, BEL-average, and BEL-optimal algorithms, respectively. All plots inclucde underlying contour lines representing the level sets of the potential.
@@ 4.2 Shape Processes
Our findings are given in Table 1 and show that BEL average outperforms competing methods. We plot a trajectory from the conditioned process learned via BEL average, alongside a trajectory from the unconditioned process in Figure 8. The next best method is Heng et al. (2022), which can be interpreted as first applying BEL to the time-reversed process, approximating the transition densities via Gaussians (24a), and using a fixed start point instead of amortisation. However, since their approach involves time-reversal, they do not take gradients of the drift or diffusion terms. BEL average not only uses this gradient information, but non-local information about the score.
@@ 4.3 Image Experiments
Next, we condition the resulting SDE to produce images with a specified upper-left corner. Importantly, this conditioning does not require adding artificial noise to the observation. More specifically, the law of $X_{T}$ is by design the data distribution. We condition on $Y=G(X_{T})$ where G selects the upper left corner of the image. Once trained we can therefore choose an arbitrary upper left corner $y$ and sample from images matching this, using the learned additional drift $u(X_{t},y)$. Note we only need train once and then can sample using any corner. We demonstrate the result by conditioning on the upper-left corner of a jacket image in Figure 6.
@@ 5 Conclusion
In this work, we introduced a novel class of loss functions to estimate the control of a conditioned diffusion process. When applied to diffusion models, we discover novel methods to add controls to an already trained diffusion model. As a byproduct, we also find novel denoising score matching objectives, which can be used for time reversal of SDEs, or to train a diffusion model. Our approach leverages Malliavin calculus and integration by parts to handle singular losses, enabling a more stable and accurate estimation. Thanks to its generality, the framework accommodates a wide range of conditioning types—continuous, discrete, noisy, or noiseless—and and could be extended to manifold-valued and infinite-dimensional settings in future works.
@@ Acknowledgments
EB is supported by the Novo Nordisk Foundation grant NNF18OC0052000, a research grant (VIL40582) from VILLUM FONDEN, and UCPH Data+ Strategy 2023 funds for interdisciplinary research. GD and JP were supported by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
@@ Appendix A Adjoint SDE for Calculating $\mathcal{S}_{s}$
making use of the semigroup property $J_{t\mid s}=J_{t\mid s+\delta t}J_{s+\delta t\mid s}$.
@@
where $\tau$ can be any number in $\tau\in[s,s+\delta t]$. For $\tau=s$, we have $J_{s\mid s}=\text{Id}$. The term $J_{s+\delta t\mid s}$ can be approximated using an Euler-Maryuama step on 7, leading to $J_{s+\delta t\mid s}\approx\text{Id}+\delta t\nabla b_{s}(X_{s})+\nabla\sigma_{s}(X_{s})\delta B_{s}$.
@@ Appendix B Denoising Score Matching and Tweedies Formula
Here we show how our results can be applied to derive a formula for the score $\nabla\log p_{t}(x)$ of a diffusion process. This could be done by applying Proposition 2.4 to a trivial conditioning $Y=0$ and treating the reverse dynamics of the SDE (6). However, it is instructive to rederive the analogous equation in this simplified setting, which we do in this section. We will see that it generalises Tweedie’s formula.
@@
and denote by $p_{t}$ the density of $X_{t}$. Then we have the following representation and score matching objective:
@@
$\nabla\log p_{t}(x)=\mathbb{E}\left[\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t\mid s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\mid X_{t}=x\right].$
Here $\alpha_{s}^{\prime}$ is a function which satisfies $\int_{0}^{t}\alpha_{s}^{\prime}\,\mathrm{d}s=1$, and $J_{t\mid s}$ is the Jacobian which follows the flow (7).
@@
$\mathcal{L}(u)=\mathbb{E}\left[\|u_{t}(X_{t})-\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t\mid s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\|^{2}\right]$ (39)
has a unique minimiser given by $u_{t}(x)=\nabla\log p_{t}(x)$.
Remark B.2.
The score matching loss (39) is a generalisation of (42) below to nonlinear SDEs. Note that in contrast to (42), the construction in (39) straightforwardly generalises to Riemannian manifolds by replacing the squared norm and the Brownian motion by their Riemannian counterparts (Hsu, 2002), and noting that $J_{t|s}$ maps between the tangent spaces $T_{X_{s}}M$ and $T_{X_{t}}M$ in such a way that the stochastic integral is well defined. The loss (39) also generalises to infinite dimensions, replacing the squared norm by a Hilbert space norm and the Brownian motion by an appropriate Wiener process. The relationship between the Wiener process and the chosen norm are however more intricate. In the diffusion model case (the linear case), they have been worked out in Pidstrigach et al. (2024, Section 2.4, Section 6).
@@
$D_{s}\varphi(X_{t})=\nabla\varphi(X_{t})J_{t\mid s}\sigma_{s}(X_{s}),$
@@
$\nabla\varphi(X_{t})=D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t\mid s}^{-1},$
@@
$\nabla\varphi(X_{t})=\int_{0}^{t}D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t\mid s}^{-1}\alpha_{s}^{\prime}\,\mathrm{d}s$
@@
$\mathbb{E}[\nabla\varphi(X_{t})]=\mathbb{E}\left[\int_{0}^{t}D_{s}\varphi(X_{t})\sigma_{s}(X_{s})^{-1}J_{t\mid s}^{-1}\alpha_{s}^{\prime}\,\mathrm{d}s\right]=\mathbb{E}\left[\varphi(X_{t})\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t\mid s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\right].$
@@
$=\frac{1}{p_{t}(x)}\mathbb{E}\left[\delta_{x}(X_{t})\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t\mid s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\right]=\mathbb{E}\left[\int_{0}^{t}(\sigma_{s}(X_{s})^{-1}J_{t\mid s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\mid X_{t}=x\right].$
@@
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}\left[X_{t}-e^{-t/2}X_{0}\mid X_{t}=x\right],$ (41)
and, in particular, the loss (39) simplifies to the well-known denoising score matching loss:
@@
$\nabla\log p_{t}(x)=\mathbb{E}\left[\int_{0}^{t}(\sigma(X_{s})^{-1}J_{t\mid s}^{-1})^{\top}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\mid X_{t}=x\right]=\mathbb{E}\left[\int_{0}^{t}e^{\frac{t-s}{2}}\alpha_{s}^{\prime}\,\mathrm{d}B_{s}\mid X_{t}=x\right].$
@@
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}\left[\int_{0}^{t}e^{-\frac{t-s}{2}}\mathrm{d}B_{s}\mid X_{t}=x\right].$ (43)
@@
$\nabla\log p_{t}(x)=\frac{1}{1-e^{-t}}\mathbb{E}[X_{t}-e^{-t/2}X_{0}\mid X_{t}=x].$
@@ C.1.3 Results
Here we provide the tables with the results for our controlled environment experiments for Brownian motion Table 4.
@@
Table 4: Performance of various algorithms for conditioning a Brownian motion (see Section 4.1.2).
Table 5: Dimension 1
Loss MV Dist
BEL average $8.1\times 10^{-1}$ $1.6\times 10^{-1}$
BEL first $2.9\times 10^{-1}$ $8.4\times 10^{-2}$
BEL last $1.1\times 10^{0}$ $1.1\times 10^{0}$
BEL optimal $8.3\times 10^{-1}$ $1.9\times 10^{-1}$
Reparametrization Trick $9.2\times 10^{-1}$ $2.1\times 10^{-1}$
Table 6: Dimension 10
Loss MV Dist
BEL average $13.4\times 10^{-}$ 1 $13.0\times 10^{-}$ 1
BEL first $12.1\times 10^{-}$ 1 $13.3\times 10^{-}$ 1
BEL last $15.9\times 10^{-}$ 1 $1.1\times 10^{0}$
BEL optimal $13.1\times 10^{-}$ 1 $13.0\times 10^{-}$ 1
Reparametrization Trick $15.5\times 10^{-}$ 1 $15.2\times 10^{-}$ 1
@@ C.2 Shape Processes
Figure 8: We plot one trajectory from the unconditioned stochastic shape process in (a), started at a circle of radius 1 and one trajectory from the process conditioned via the proposed method BEL average to end at a circle of radius 1.5 in (b).
@@ Appendix D Proofs
In order to prove Proposition 2.4 (and later Lemma 3.3), we now prove that the score is a martingale, when conditioned on the observation $Y$:
Lemma D.2.
Let $s\geqslant t$, then we have that
$\nabla_{X_{t}}\log p(Y\mid X_{t})=\mathbb{E}[\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\mid Y,X_{t}].$ (48)
Proof.
We have
$\nabla_{x_{t}}\log p(Y\mid X_{t})$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int p(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{t})\mathrm{d}x_{T}$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int p(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{s}=x_{s})p(X_{s}=x_{s}\mid X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int p(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{s}=x_{s})p(X_{s}=x_{s}\mid X_{t})\nabla_{X_{t}}\log p(X_{s}=x_{s}\mid X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$\qquad=\mathbb{E}[\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\mid X_{t},Y].$
@@
$\nabla\log p_{T\mid s}(X_{T}=x_{T}\mid X_{s}=x_{s})$ $=~{}\frac{1}{p_{T\mid s}(X_{T}=x_{T}\mid X_{s}=x_{s})}\nabla_{x_{s}}\mathbb{E}[\delta_{x_{T}}(X_{T})\mid X_{s}=x_{s}]$
$=~{}\mathbb{E}\left[\delta_{x_{T}}(X_{T})\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid X_{s}=x_{s}\right]\frac{A_{T\mid s}^{-1}}{p_{T\mid s}(X_{T}=x_{T}\mid X_{s}=x_{s})}$
@@
The final result of Proposition 2.4 follows transposing $\alpha$ and transposing the last equation; this is only a cosmetic change so that the result is more in line with algorithmic implementations.
Now, using Lemma D.2 for $t=T$ we get that
$\nabla\log p(Y=y\mid X_{s}=x_{s})$ $=\mathbb{E}[\nabla_{X_{s}}\log p_{T\mid s}(X_{T}\mid X_{s})\mid Y,X_{s}]$
$=\mathbb{E}\left[\mathbb{E}\left[\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid X_{s},X_{T}\right]\mid Y,X_{s}\right]$
$=\mathbb{E}\left[\int_{s}^{T}(\sigma_{t}(X_{t})^{-1}J_{t\mid s}\alpha_{t}^{\prime})^{\top}\mathrm{d}B_{t}\mid Y,X_{s}\right].$